OpenStudy (anonymous):

5. What is the approximate radius of a sphere whose volume is 1,349 cm³? (1 point) 5.7 cm 6.9 cm 11 cm 14.7 cm

OpenStudy (anonymous):

well if we know that Vs=4/3pir^3 then r=(Vs/1.33pi)^1/3

OpenStudy (campbell_st):

Well the formula is \[V = \frac{4}{3} \pi r^3\] substitute your volume \[1349 = \frac{4}{3} \pi r^3\] now solve for r...

OpenStudy (goformit100):

don't use calculator

OpenStudy (anonymous):

omg this is soo hard -_-

OpenStudy (campbell_st):

start by multiplying both sides by 3 to eliminate the denominator of 3

OpenStudy (goformit100):

well said @campbell_st

OpenStudy (anonymous):

ok.....

OpenStudy (anonymous):

can someone just me give me the answer???!!!!!!!!!!!!!!

OpenStudy (goformit100):

Whats ur answer ?

OpenStudy (anonymous):

i dont have one that is why im asking.....

OpenStudy (campbell_st):

wellI I'll help you calculate it... but I won't give you the answer...

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

what did you get when you multiplied both sides by 3

OpenStudy (anonymous):

formula for volume is \[V =4\pi r^3/3\] then put the value of V in given formula that is \[1349=4 \pi r^3/3\] \[1349*3/(4*\pi)=r^3\] then you will get your result.

OpenStudy (goformit100):

GREAT @ashu97 :)

OpenStudy (anonymous):

thank s dear..